In clinical tests of Lipitor,  270 subjects were given a placebo and 7 had allergic reactions. Among the 863 subjects treated with the drug, 8 had allergic reactions. Construct two 95% CI's for the proportion of allergic reactions and compare the results.
Answer #1

Solution :

Given that,

Point estimate = sample proportion = = x / n = 270 / 863 = 0.313

1 - = 1 - 0.313 = 0.687

Z/2 = 1.96

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 1.96 * (((0.313 * 0.687) / 863)

Margin of error = E = 0.031

A 95% confidence interval for population proportion p is ,

- E < p < + E

0.313 - 0.031 < p < 0.313 + 0.031

0.282 < p < 0.344

The 95% confidence interval for the population proportion p is : 0.282 , 0.344